OpenStudy (anonymous):

Evaluate the function at the given value of the variable. f(x) = √4 – 4(x) ; x = - 4, 11

OpenStudy (anonymous):

For f(x)= √ 4-4(x); x=-4,11 f(-4)= √ 4-4(-4) f(-4)= √ 4+16 f(-4)= √ 20 Will substitute 11 for x f(x)= √ 4-4 f(11)= √ 4-4(11) (this becomes 44-4) f(11)= √ -40 since -40 is a negative number, then f(11) is not a real number. Do I have it correct?